# CardioLive: Empowering Video Streaming with Online Cardiac Monitoring

### v1

@@ 3.2.1 Video Branch Design
Temporal Differential Block (TDB): The input video frames $\Pi$ will first be processes as, i.e.,
$\dot{\Pi}_{i,j}^{t}=\Pi_{i,j}^{t}-\Pi_{i,j}^{t-1}.$ (3)
Note that in online learning, we only have past information, so we perform backward differentiation. The key idea is, we treat the psychological activities as tiny local ”motions”. It efficiently captures the changes between consecutive frames [60]. Furthermore, TDB plays a crucial role in isolating dynamic features while suppressing static components present in the video data, as stated in Eq. (1). Furthermore, temporal difference enhances the contrast of the cardiac signal $h(t)$ within the latent space, facilitating more effective feature extraction and subsequent analysis. Thereafter, they are fed into convolution networks and upsampled to meet the length of video features. It is also imperative to capture the static information inherent in the video frames. To this end, we integrate a parallel pathway to process the original video frames, allowing for a more comprehensive understanding of the environment. We then introduce lateral connections to facilitate fusion of static and dynamic information.
Static and Dynamic Path: In addition to dynamic modeling through the temporal difference block described earlier, it is imperative to capture the static information inherent in the video frames. While the temporal difference mechanism offers a means to extract low-resolution and sparse representations, it inevitably overlooks crucial static details such as appearance and illumination. Recognizing the significance of this static information, we integrate a parallel pathway to process the original video frames. This parallel pathway begins with passing the original video frames through a convolution block, followed by downsampling using max pooling , i.e.,
$\mathcal{S}(\bm{\Pi^{t}})=\operatorname{Downsample}(\operatorname{CNN}(\bm{{\Pi}^{t}})).$ (4)
This can preserve static information, allowing for a more comprehensive understanding of the environment. To ensure mutual awareness and coherence between the dynamic and static pathways, we introduce lateral connections to facilitate fusion. Specifically, bidirectional connections are employed to merge features by downsampling the static branch and upsampling the dynamic branch , i.e.,
$\hat{\bm{\Pi}}=\cdot\mathcal{D}(\bm{\Pi})\oplus\cdot\mathcal{S}(\bm{\Pi})$ (5)
This fusion mechanism enables the model to effectively leverage insights gained from both temporal dynamics and static features, enhancing overall performance and interoperability.
@@
Motion-Aware Aggregation (MAA): The above design incorporates temporal information with static and dynamic modeling. After lateral fusion, we pass the intermediate latent to the bottleneck block , which gradually reduces the channels of the feature while being effective to extract the spatial information and increase the expressive power. We recognize the importance of spatial modeling in mitigating the motion noise from head movement. Unlike video recognition tasks, where the relative location of the pixel is vital, we care more about how to track the variations of these pixels over time. To this end, we introduce a self-attention mechanism for frame-wise aggregation between consecutive frames. Our goal is to establish a mapping between temporal pixel variations and consecutive spatial information. Given the latent space $\bm{\hat{\Pi}}\in\mathbb{R}^{\hat{T}\times\hat{C}\times\hat{H}\times\hat{W}}$, we query the one pixel at time $t$, i.e., $\hat{\Pi}^{t}_{i,j}$ and compute the attention with previous frame,
@@
This mechanism scrutinizes pixel displacements across consecutive frames, akin to tracing the path of movement within a sequence of images. Each pixel’s attention weight encapsulates its significance in depicting motion, allowing the model to recognize subtle shifts and fluctuations over time.
This mechanism allows the model to selectively attend to relevant spatial features and enhance the temporal motion modeling. In essence, this scheme can be conceptualized as a motion vector, dynamically tracking and encoding the intricate temporal variations within video data. By leveraging the self-attention mechanism, The model scrutinizes pixel displacements across consecutive frames, akin to tracing the path of movement within a sequence of images. Each pixel’s attention weight encapsulates its significance in depicting motion, allowing the model to recognize subtle shifts and fluctuations over time. Through this process, the model elucidates the underlying motion patterns embedded within the video stream.
Frequency-Aware Block (FAB): After applying motion attention aggregation, we acquire the enhanced feature $\bm{\check{\Pi}}\in\mathbb{R}^{T\times\hat{C}\times\hat{H}\times\hat{W}}$. Our previous focus has been on modeling video dynamics in the temporal domain. These are very effective designs for cardiac time series learning. Moreover, given the intrinsic property of $h(t)$, which turns out to be a quasi-periodic signal, it becomes imperative to incorporate frequency features into our analysis. Here, the term ”frequency” does not merely refer to the spectrum of color space within the video; rather, we aim to capture the underlying frequency variations of pixels over time. Inspired by DTF [43], we attempt to explicitly incorporate FFT in our design. For each pixel $\bm{\check{\Pi}_{i,j}}\in\mathbb{R}^{T\times\hat{C}}$, we apply FFT along the temporal dimension to acquire the feature spectrum $\Psi_{\bm{\check{\Pi}_{i,j}}}(f)$. To capture the frequency information, we introduce a learnable frequency filter $\Psi_{G}(f)\in\mathbb{R}^{\hat{C}\times N_{f}}$. We use IFFT to get the modulated temporal feature. With FCB, we can enlarge the receptive field and profile cardiac time series with frequency constraints.
Irregular Sampled Time Embedding: Another challenge of online cardiac learning is the fluctuating FPS. To this end, we introduce the timestamp feature to handle the irregular sampled time learning. Technically, we can acquire the set of timestamps $\{t_{i}\}_{i=1}^{N_{v}}$ for each frame. We incorporate a timestamp embedding $E_{t}$ design and fuse it with $\bm{\Theta_{E_{v}}(\Pi)}$. Specifically, we employ a frequency embedding scheme, which computes triangle embedding based on a geometric progression of frequencies up to $f_{m}$. We first derive a set of frequencies with the size of embedding dimension $N_{t_{d}}$, i.e.,
@@
where $k=1,\cdots N_{t_{d}}/2$. Then the angle for each timestamp $i$ is given by
$\theta_{i}^{k}=t_{i}\cdot\omega^{k}\cdot 2\pi.$ (9)
Finally, the timestamps are embedded through trigonometric encoding by concatenating sine and cosine values for each angle.
3.2.2 Audio Branch Design
We then introduce the design for the audio encode $E_{a}$. As discussed in §3.1, human speech is modulated by the time-varying filter $R(f)$. And the cardiac series is encapsulated in $R(f)$. Inspired by this filtering process, we opt to emulate it within our design. We target directly processing raw audio in our case. We will justify the rationale first, followed by our design.
Raw Audio: Traditional audio-based learning often leverages mel-spectrogram, a common practice for tasks like speech recognition. However, this method may not be suitable for our task. Our predictions, $h(t)$, manifest as quasi-periodic signals, ideally shown as straight lines on a mel-spectrum. But because cardiac activities are variable, these lines will exhibit randomness on a temporal-frequency map. Also, the location of the ”straight” line has physical meanings, rather than a simple pattern. Therefore, we resort to learning from the raw audio signals directly. The key insight is, the process of producing speed from our vocal organs is composed of several acoustic filters, as indicated in §3.1. We can simulate the effect of filters and incorporate them in our design.
Temporal-Frequency Filter (TFF): The temporal format of Eq. (2) can be rewritten as $\xi(t)=l(t)\ast r(t)$, where $\xi(t)$ is the speech signal. $l(t)$ represents the source of the sound while $r(t)$ is combination of source filters. To this end, we adopt the SincNet [49], which can be expressed as,
@@ 3.2.2 Audio Branch Design
$f_{i,2}^{\theta}$ and $f_{i,1}^{\theta}$ denotes the two cutoff frequencies. We can treat the two cutoff frequencies as learnable parameters. We then perform convolution between $r_{i}(t)$ and raw audio $\xi(t)$. They will be fed into 1D convolution blocks for feature extraction.
3.2.3 Fusion Block Design
Until now, we have handled the video feature $\Theta_{E_{v}}(\Pi)$ and audio feature $\Theta_{E_{a}}(\Xi)$. We now present the design of the fusion network. We opt for the late-fusion scheme, as the relationships between audio and cardiac activity, as well as video and cardiac activity, are not initially apparent. Within the fusion block, we aim to address two challenges: 1) aligning the audio and video features along the temporal domain, and 2) handling the sampling rate mismatch between the audio and video features. To do so, we propose a multi-head temporal attention fusion block. Subsequently, the fused feature will be passed through linear fully connected layers. To achieve this, we adopt a temporal attention-based fusion scheme. Technically, we exploit video features as the query, and audio features as the key and value , i.e.,
@@ 3.2.3 Fusion Block Design
3.2.4 Loss
In this part, we will elaborate our loss function design.
In the time domain, we calculate the focal loss:
$\mathcal{L}_{\text{dis}}=\frac{1}{N}\sum_{i=1}^{N}\left(\epsilon\cdot\left(1-\operatorname{e}^{-\mathcal{M}\left(\hat{h(t_{i})},h(t_{i})\right)}\right)^{\eta}\cdot\mathcal{M}\left(\hat{h(t_{i})},h(t_{i})\right)\right),$ (12)
where $\hat{h(t)}$ is the predicted cardiac series. $\epsilon$ and $\eta$ are the balancing coefficient and focusing coefficient, respectively. $\mathcal{M}(\cdot,\cdot)$ represents the mean square error (MSE). Given the inherent complexity and variability of physiological signals, focal loss offers a more robust framework for capturing peaks. Besides, we also introduce similarity loss,
$\mathcal{L}_{\text{sim}}=1-\frac{\sum_{i=1}^{N}\left(\hat{h(t_{i})}\cdot h(t_{i})\right)}{\sqrt{\sum_{i=1}^{N}\hat{h(t_{i})^{2}}}\cdot\sqrt{\sum_{i=1}^{N}h(t_{i})^{2}}}.$ (13)
The range of $\mathcal{L}_{\text{sim}}$ is $[0,2]$, where $\mathcal{L}_{\text{sim}}=0$ represents perfect alignment and is the optimization goal. Additionally, as we are learning quasi-periodic signal, we incorporate spectral loss as well. Specifically, we calculate the FFT-MSE loss,
$\mathcal{L}_{\text{freq}}=\mathcal{M}\left(\Psi_{\hat{h(t)}}(f),\Psi_{h(t)}(f)\right),$ (14)
where $\Psi(\cdot)$ represents the spectrum function. To wrap up, we calculate the weighted sum of the above three losses,
$\mathcal{L}_{\text{all}}=\alpha\cdot\mathcal{L}_{\text{dis}}+\beta\cdot\mathcal{L}_{\text{sim}}+\gamma\cdot\mathcal{L}_{\text{freq}},$ (15)
where $\alpha$, $\beta$ and $\gamma$ are weights to balance the loss items. Thus far, we present the design of CardioNet. To our best knowledge, it is the first audio-visual network designed for cardiac learning. We will demonstrate the performance in §5.
4 CardioLive Design
In this section, we will introduce the design of CardioLive. We will introduce the design goal of CardioLive in §4.1, followed by our detailed designs of service in §4.2 and §4.3. We will introduce the preprocessing in §4.4.
@@ 4 CardioLive Design
4.1 Design Goal
Modern video streaming systems are complicated, and integrating OCM into them is non-trivial. As shown in Fig. 7, the content is sent through cloud servers spanning across different locations globally. Besides running the data center and cloud computing, these video streaming systems offer a range of application services, such as content summarization, transcriptions, and AI-driven interactive features. Note that for content providers like YouTube, Netflix, and many VoD providers, integrating new features is relatively straightforward because they can preload resources in their data centers. However, this does work well with streaming systems with live content generation and interactions. On the other hand, deploying cardiac monitoring on edge devices is also valuable. Users will be concerned about how the sensitive data are communicated over the network.
Therefore, to achieve SoD cardiac monitoring service, we need to both consider deploying the cardiac monitoring services on the edge ends, e.g., browsers, and application services. Notably, direct access to data that manufacturers possess is often restricted by stringent privacy regulations affecting external developers. To this end, we aim to package CardioLive into a service, which both end users and manufacturers can readily access. At a high level, we are not concerned about specific implementations on specific platforms, but aim to develop CardioLive as a microservice. We utilize data hooks to capture video and audio streams, organizing them into buffer queues as data packets, which will be fed to the inference engines, as detailed below.
4.2 Buffer Design
Data Hook: We will first introduce data hooks to get the video and audio streams, namely onVideoDataReceived() and onAudioDataReceived(). Meeting platforms like Zoom usually support in-app bots that virtually participate in calls. We can leverage the bots to access the raw data streams, as shown in Fig. 8(a). Meanwhile, increasingly more video streaming systems are based on web pages, e.g., YouTube, Bilibili, etc. Direct accessing the video streams of this platform is rather complicated and violates the policies. To this end, we leverage WebGL and WebAudio that exist in modern browsers to get the data streams, as shown in Fig. 8(b). The browsers usually provide the Document Object Model (DOM), a programming interface to manipulate the structure, style, and content of web content. Our service will first access the canvas, an element for graphics on a web page, through DOM. The canvas offers a bitmap where each pixel can be individually manipulated. We get the rendering context through WebGL, which operates as a rendering context of canvas using the underlying GPU. We create an offscreen canvas that is rendered off the main thread and read the pixels through WebGL, preventing it from interfering with the normal UI updates. Meanwhile, we capture the audio from the video element through WebAudio, a versatile framework to handle audio operations on the web. We record the timestamp of the audio and video as well. Through the data hook, we can acquire the video and audio streams. Then we will construct them into data packets and buffer queues.
@@ 4.2 Buffer Design
Data Packet: Normally, audio and video are encoded in separate ways. In meeting platforms, the video frames are usually encoded in YUV format. They are designed for the best transmission efficiency. Encoding the data in YUV space allows fewer total bits of space in a video stream for the colors to be shown. To recover the original RGB streams, we have first to decode the YUV streams. To reduce the cost of decoding, we adapt a streaming-based decoding pipeline from GStreamer [3]. We set the appsink property for receiving the RGB data and assign appsrc for handling YUV encoding. We set the transformation in asynchronous mode so that the incoming frames will not conflict with the current operations. After that, we will construct the collected frames in buffers.
Then we feed the video-audio pair into the forwarding packets. For audio and video streams, we apply the same packet format, which contains a unique header, an identifier, the data size, timestamps, and the encoded payload data, as illustrated in Fig. 9. The unique header is designed to judge whether the packet is correctly constructed and not mixed with other packets. The identifier is assigned to indicate whether it is audio or video data packets. We embed the received timestamps to denote the sequence of the video and audio, which will be further used for synchronization.
@@
4.3 Service Design
We abstract our system as a plug-and-play service. Our service first gets the hooked video and audio packets as the input. The data will be preprocessed and fed to the inference engine for output. Our design overcomes the two challenges: fluctuating FPS and unsynchronized audio and video streams.
Changing FPS: The fluctuating FPS will lead to two subproblems. Initially, the video streaming systems will ideally have 30 FPS but in reality undersampled at the receiver’s end, as illustrated in Fig. 10, with some outliers present as well. Additionally, the frame rate is not constant, resulting in a varying number of frames within a given window. However, our model assumes a fixed 4-s input, with 120 frames of video (30 FPS) and 32000 samples of audio (8kHz). In other word, we have to adapt the real input size to the model. To this end, instead of padding empty frames at the end, we duplicate one single frame circularly. For instance, if the actual FPS is 25, we insert an additional identical frame after every 5 frames to approximate a smoother transition to 30 FPS. Any remaining gaps at the end of the sequence are filled by repeating the last frame. As for overlarge FPS, we downsample the frames. For the audio clips, as 8kHz is much lower than the typical sampling rates (usually 32kHz or 44.1kHz) in modern video streaming systems, we can concatenate the received audio chunks and safely downsample them to 8kHz.
@@ 4.3 Service Design
Audio Video Synchronization: The audio-visual misalignment is a more severe issue than changing FPS. As we are hooking audio and video from separate channels, they are likely to lose synchronization with the increase of time. As can be seen from Fig. 8(b), the starting time of the audio and video will be misaligned quickly with accumulating drifts. To overcome this issue, we develop a scheme to ensure the audio and video chunks are synchronized before sending to the inference engine. Given the audio and video streams $S_{a}(t)$ and $S_{b}(t)$, they will be extended to the buffer queues $Q_{a}(t)$ and $Q_{v}(t)$, respectively. We also maintain $t_{a}$ and $t_{v}$ as the starting time of audio and video chunks, respectively. We denote $\Delta t_{n}=t_{a}^{k}-t_{v}^{k}$ as the temporal drift between audio and video streams at $k$-th trial. To mitigating the continuously increasing $\Delta t_{n}$, we align the start time at each step $k$ as,
$t_{\text{start}}^{k}=\max(t_{a}^{k},t_{v}^{k}),$ (16)
when $\Delta t_{n}$ is larger than the threshold $\epsilon_{t}$. We use $\epsilon_{t}=0.3$s. Then the ending time will be determined by $t_{\text{end}}^{k}=t_{\text{start}}^{k}+t_{w}$, where $t_{w}$ is the window lengths. Note that we are adopting a sliding window scheme, with window length $t_{w}$ and step length $t_{s}$. For the next window, the start time will be updated by finding the timestamp closest to, i.e., $t_{a}^{k+1}=t_{a}^{k}+t_{s}$ and $t_{v}^{k+1}=t_{v}^{k}+t_{s}$.
@@
Meanwhile, we will pop the items that have been processed from the buffer queues, i.e.,
$Q_{a}(t)$ $=Q_{a}(t)\textbackslash\{S_{a}(t)|t<t_{a}^{k+1}\}$ (17)
$Q_{v}(t)$ $=Q_{v}(t)\textbackslash\{S_{v}(t)|t<t_{v}^{k+1}\}$
We then feed the synchronized pairs for inference.
4.4 Preprocessing
In this section, we will discuss the preprocessing pipelines. We use OpenCV face detector to find the faces. We also perform voice activity detection to segment the talking period. Additionally, we need to separate multiple persons, if any, and match their audio and videos.
Audio-Visual Matching: To realize the matching between speaking clips and facial hints, we adopt a cross-attention scheme [56, 32]. Specifically, after the encoders, we get two features $M_{a}$ and $M_{v}$. These features are expected to encapsulate relevant speaking activities by employing temporal encoders [31, 15]. To fuse the audio and video features, The audio features $M_{a}$ are integrated with the video data by treating $M_{v}$ as the target for querying through an attention framework. Conversely, the video features $M_{v}$ interact with $Q_{a}$, representing the audio query sequences. The outputs are concatenated together along the temporal direction.
5 Evaluation
In this section, we systematically evaluate CardioLive. We aim to answer the following questions:
  1. 1.
    To what extent does CardioLive outperform existing video-only and audio-only solutions?
  2. 2.
    How effectively does CardioLive perform across various scenarios, including edge cases?
  3. 3.
    How does the streaming service of CardioLive perform when integrated into real-world video streaming systems?
We perform comparison studies with the state-of-the-art (SOTA) video-based solutions and audio-based solutions. We mainly leverage our self-collected dataset. We use the following metrics to evaluate the accuracy of the model: Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Mean Absolute Percentage Error (MAPE).

### v2

@@ III-B1 Video Branch Design
Temporal Differential Block (TDB): The input video frames $\Pi$ will first be processed as, i.e., $\dot{\Pi}_{i,j}^{t}=\Pi_{i,j}^{t}-\Pi_{i,j}^{t-1}.$ Since we only have past information, we perform backward differentiation. The key idea is, we treat the psychological activities as tiny local ”motions”. It efficiently captures the changes between consecutive frames [wang2021tdn]. Furthermore, TDB plays a crucial role in isolating dynamic features while suppressing static components present in the video data. Temporal difference enhances the contrast of the cardiac signal $h(t)$ within the latent space, facilitating more effective feature extraction and subsequent analysis. Thereafter, they are fed into convolution networks and upsampled to meet the length of video features. It is also imperative to capture the static information inherent in the video frames. To this end, we integrate a parallel pathway to process the original video frames, allowing for a more comprehensive understanding of the environment. We then introduce lateral connections to facilitate fusion of static and dynamic information.
@@
Motion-Aware Aggregation (MAA): After lateral fusion, we pass the intermediate latent to the bottleneck block to extract the spatial information and increase the expressive power. We recognize the importance of spatial modeling in mitigating the motion noise from head movement. Unlike video recognition tasks, where the relative location of the pixel is vital, we care more about how to track the variations of these pixels over time. To this end, we introduce a self-attention mechanism for frame-wise aggregation between consecutive frames. Our goal is to establish a mapping between temporal pixel variations and consecutive spatial information. Given the latent space $^{\hat{\Pi}}$, we query the one pixel at time $t$, i.e., $\hat{\Pi}^{t}_{i,j}$ and compute the attention with previous frame,
@@
This mechanism scrutinizes pixel displacements across consecutive frames, akin to tracing the path of movement. Each pixel’s attention weight encapsulates its significance in depicting motion, allowing the model to recognize subtle shifts and fluctuations over time.
Frequency-Aware Block (FAB): After applying motion attention aggregation, we acquire the enhanced feature $\check{\Pi}$}. Our previous focus has been on modeling video dynamics in the temporal domain. These are very effective designs for cardiac time series learning. Moreover, given the intrinsic property of $h(t)$, which turns out to be a quasi-periodic signal, it becomes imperative to incorporate frequency features into our analysis. Here, the term ”frequency” does not merely refer to the spectrum of color space within the video; rather, we aim to capture the underlying frequency variations of pixels over time. Inspired by DTF [long2022dynamic], we attempt to explicitly incorporate FFT in our design. For each pixel $^{\check{\Pi}_{i,j}}\in\mathbb{R}^{T\times\hat{C}}$, we apply FFT along the temporal dimension to acquire the feature spectrum $\Psi_{{\check{\Pi}_{i,j}}}(f)$. To capture the frequency information, we introduce a learnable frequency filter $\Psi_{G}(f)\in\mathbb{R}^{\hat{C}\times N_{f}}$. We use IFFT to get the modulated temporal feature. With FCB, we can enlarge the receptive field and profile cardiac time series with frequency constraints.
Irregular Sampled Time Embedding: Another challenge is the fluctuating FPS. To this end, we add the timestamp feature to handle the irregular sampled time. Given the set of timestamps $\{t_{i}\}_{i=1}^{N_{v}}$, we design the timestamp embedding $E_{t}$ and fuse it with $\Theta_{E_{v}}(\Pi)$}. Specifically, we employ a frequency embedding scheme, which computes triangle embedding based on a geometric progression of frequencies up to $f_{m}$. We first derive a set of frequencies with the size of embedding dimension $N_{t_{d}}$, i.e.,
@@
where $k=1,\cdots N_{t_{d}}/2$. Then the angle for each timestamp $i$ is given by $\theta_{i}^{k}=t_{i}\cdot\omega^{k}\cdot 2\pi.$ Finally, the timestamps are embedded through trigonometric positional encoding.
III-B2 Audio Branch Design
We then introduce the audio encoder.
Raw Audio: Traditional audio-based learning often leverages mel-spectrogram. However, this method may not be suitable for our task. Our predictions, $h(t)$, manifest as quasi-periodic signals, ideally shown as straight lines on a mel-spectrum. But because cardiac activities are variable, these lines will exhibit randomness on a temporal-frequency map. Also, the location of the ”straight” line has physical meanings, rather than a simple pattern. Therefore, we resort to learning from the raw audio signals directly. The key insight is, the process of producing speed from our vocal organs is composed of several acoustic filters, as indicated in §III-A. We can simulate the effect of filters and incorporate them in our design.
Temporal-Frequency Filter (TFF): The cardiac effect on the speech can be seen as a match filter. To this end, we adopt the SincNet [ravanelli2018speaker], which can be expressed as,
@@ III-B2 Audio Branch Design
$f_{i,2}^{\theta}$ and $f_{i,1}^{\theta}$ denotes the two cutoff frequencies. We can treat the two cutoff frequencies as learnable parameters. We then perform a convolution between $r_{i}(t)$ and the raw audio $\xi(t)$. They will be fed into 1D convolution blocks for feature extraction.
III-B3 Fusion Block Design
We now present the design of the fusion network. We opt for the late-fusion scheme, as the relationships between audio and cardiac activity, as well as video and cardiac activity, are not initially apparent. Moreover, late fusion provides architectural flexibility. In scenarios where audio is absent (such as user silence), our system gracefully defaults to video-only inference while maintaining consistency. Within the fusion block, we aim to address two challenges: 1) aligning the audio and video features along the temporal domain, and 2) handling the sampling rate mismatch between the audio and video features. To do so, we propose a multi-head temporal attention fusion block. Subsequently, the fused feature will be passed through linear fully connected layers. Technically, we exploit video features as the query, and audio features as the key and value , i.e.,
@@ III-B3 Fusion Block Design
III-B4 Loss
In this part, we will elaborate our loss function design. We include three types of loss functions, i.e., focal loss, frequency loss and similarity loss, $\mathcal{L}_{\text{all}}=\alpha\cdot\mathcal{L}_{\text{dis}}+\beta\cdot\mathcal{L}_{\text{sim}}+\gamma\cdot\mathcal{L}_{\text{freq}},$ where $\alpha$, $\beta$ and $\gamma$ are weights to balance the loss items. The focal loss $\mathcal{L}_{\text{dis}}$ excels at keeping peaks in the physiological signals [ross2017focal]. The similarity loss $\mathcal{L}_{\text{sim}}$ represents the extent of alignment. Additionally, as we are learning a quasi-periodic signal, we incorporate spectral loss $\mathcal{L}_{\text{freq}}$ as well by calculating the MSE of FFT.
IV System Design
@@
IV-A Design Goal
Modern video streaming systems are complicated, and integrating OCM into them is non-trivial. As shown in Fig. 8, the content is sent through cloud servers spanning across different locations globally. Besides running the data center and cloud computing, these video streaming systems offer a range of application services, such as content summarization, transcriptions, and AI-driven interactive features. For VoD providers, integrating new features is straightforward because they can preload resources in their data centers. However, this does work well with streaming systems with live content interactions. Meanwhile, deploying cardiac monitoring on end devices is also valuable. Users will be concerned about how the sensitive data are communicated over the network. To achieve SoD cardiac monitoring, we consider deploying it both on the ends and on the cloud. We package CardioLive into a service, which both end users and manufacturers can readily access. In other word, we are not concerned about implementations on specific platforms, yet develop CardioLive as a microservice. We elaborate on it below.
IV-B Buffer Design
Data Hook: We design data hooks to get the video and audio streams, namely onVideoDataReceived() and onAudioDataReceived(). Meeting platforms like Zoom usually support internal bots that join the calls. We can leverage the bots to access the raw data streams, as shown in Fig. 9(a). Meanwhile, most of the video streaming systems are based on web pages, e.g., YouTube, Bilibili, etc. Directly accessing the video streams of this platform is rather complicated and violates the policies. To this end, we leverage WebGL and WebAudio that exist in modern browsers to get the data streams, as shown in Fig. 9(b). The browsers usually provide the Document Object Model (DOM), a programming interface to manipulate the structure, style, and content of web content. Our service will first access the canvas, an element for graphics on a web page, through the DOM. The canvas offers a bitmap where each pixel can be individually manipulated. We get the rendering context through WebGL and create an offscreen canvas that is rendered off the main thread and read the pixels through WebGL, preventing it from interfering with the normal UI updates. Meanwhile, we capture the audio from the video element through WebAudio, a versatile framework to handle audio operations on the web. We record the timestamp of the audio and video as well. Through the data hook, we can acquire the video and audio streams. Then we will construct them into data packets and buffer queues.
Data Packet: Normally, audio and video are encoded in separate ways. In meeting platforms, the video frames are usually encoded in YUV format, designed for the best transmission efficiency. To recover the original RGB streams and reduce the cost of decoding, we adapt a streaming-based decoding pipeline from GStreamer [GStreamer]. We set the appsink property for receiving the RGB data and assign appsrc for handling YUV encoding. The transformation is in asynchronous mode so that the incoming frames will not conflict with the current operations. After that, we construct the collected frames in buffers. We feed the video-audio pair into the forwarding packets. For audio and video streams, we apply the same packet format, which contains a unique header, an identifier, the data size, timestamps, and the encoded payload data, as illustrated in Fig. 10. The unique header is designed to judge whether the packet is correctly constructed and not mixed with other packets. The identifier is assigned to indicate the audio or video data packets. We embed the received timestamps to denote the sequence of the video and audio, which will be further used for synchronization.
@@ IV-B Buffer Design
IV-C Service Design
We abstract our system as a plug-and-play service. Our service first gets the hooked video and audio packets as the input. The data will be fed to the inference engine for output. We observe and tackle the two challenges: fluctuating FPS and unsynchronized streams.
Drifting FPS: The fluctuating FPS will lead to two subproblems. Initially, the video streaming systems will ideally have 30 FPS but in reality undersampled at the receiver’s end, as illustrated in Fig. 11, with some outliers present as well. Additionally, the frame rate is not constant, resulting in a varying number of frames within a given window. However, our model assumes a fixed 4-s input, with 120 frames of video (30 FPS) and 32000 samples of audio (8kHz). In other words, we have to adapt the real input size to the model. To this end, instead of padding empty frames at the end, we duplicate a single frame circularly. For instance, if the actual FPS is 25, we insert an additional identical frame after every 5 frames to approximate a smoother transition to 30 FPS. Any remaining gaps at the end of the sequence are filled by repeating the last frame. As for overlarge FPS, we downsample the frames. For the audio clips, as 8kHz is much lower than the typical sampling rates (usually 32kHz or 44.1kHz) in modern video streaming systems, we can concatenate the received audio chunks and safely downsample them to 8kHz.